# A short-baseline neutrino beam at CERN for high-precision cross-section measurements

### v1

@@ 7 Neutrino cross-section measurements
- 7.8 Summary of the SBN@CERN physics potential
@@ 1 A high precision neutrino beam at CERN
In 2023, the convergence of the ENUBET and NuTag approaches revealed the potential to achieve neutrino-nucleus cross-section measurements with a precision comparable to electron-nucleus cross-sections. Recognizing this opportunity, the two collaborations joined forces to explore the feasibility of implementing such a facility at CERN [19]. The beamline design was optimized within the framework of Physics Beyond Colliders (PBC) at CERN, culminating in the design presented in this document. This document outlines the results of these efforts and provides the foundation for the proposal of a new high-precision, short-baseline neutrino beam at CERN, tentatively named SBN@CERN. This facility aims to deliver neutrino cross-section measurements at the percent level in the energy range relevant for DUNE and Hyper-Kamiokande. While alternative implementations of this innovative technology outside CERN are under consideration, CERN’s unparalleled infrastructure, proton driver, and existing neutrino detectors position CERN as the most cost-effective venue for realizing this project in compliance with the CERN fixed-target program for the coming decade.
The SBN@CERN beamline is detailed in Sec. 2, with a focus on its implementation at CERN, utilizing the SPS as the proton driver. Section 3 outlines the beamline instrumentation, which monitors and tracks mesons and charged leptons, while Sec. 4 discusses the neutrino detectors that leverage the advancements of the CERN Neutrino Platform. Section 5 summarizes the technique that enables a sub-percent precision measurement of the neutrino flux using lepton monitoring and tracking. Section 6 presents the two methods by which SBN@CERN can measure the neutrino energy without relying on neutrino interactions: the Narrow Band Off-Axis technique developed by ENUBET and the SBND Collaboration at Fermilab [20, 21], which achieves a precision of approximately %, and the tagging method, which can reach the ultimate precision of 1%. The expected performance of the facility for a range of reference cross-section measurements is detailed in Sec. 7.
@@ 2 Beamline design
The SBN@CERN beamline aims at producing kaons and pions with a central momentum range of . The beamline, shown in Fig. 1, was optimized to produce (ideally) a monochromatic meson beam that remains parallel throughout a 40 m-long decay tunnel. The neutrino detectors that are under consideration are discussed in Sec. 4. In the following, it is assumed to be a liquid-argon detector located downstream of the decay tunnel. It has a transverse size of m2 and a longitudinal depth of roughly (see Sec. 4.3). The meson production is achieved via a thin graphite target that is impinged by the proton beam of CERN’s SPS.
@@
At CERN, the beamline can only be effectively driven by the SPS accelerator with a proton-beam energy in the range up to . The preliminary beamline specifications are listed in Tab. 1. At PS energies, the meson yield is highly suppressed making this option not viable. If implemented at CERN, the SBN beamline will be confronted with the issue of having a limited number of protons-on-target (PoT) available since other experiments are also dependent on a high annual proton flux from the SPS. For this reason, the beamline design has been optimized to minimize the required number of PoTs.
@@ 2.1 Beamline optimization
| Particle yield | ENUBET design | Optimized SBN design |
@@ 2.2 Radiation-protection considerations
SBN@CERN is a proposed beamline featuring novel instrumentation for flux monitoring and tagging. However, radiation effects must be carefully considered, as high radiation doses may impact beamline components, silicon taggers, and beamline instrumentation. This challenge has already been addressed by the ENUBET R&D and is further mitigated by the reduction of PoT achieved through the optimization described in Sec. 2.1. Concerning the ambient dose equivalent (H*(10)) studies are currently ongoing.
@@
Although radiation damage is not critical for the SBN@CERN components up to PoT, shielding for radiation protection within the CERN complex remains an important task. The required shielding depends on the final location of SBN@CERN and is currently under evaluation. Location options are discussed in the next section.
@@ 2.3 Possible implementation at CERN
The feasibility study for a potential placement of the SBN beamline on the CERN campus is currently ongoing and is not final [35]; however, two approaches have been identified as promising and could lead to the successful implementation of the beamline.
CERN’s North Area (see Fig. 5) is usually the prime location for fixed-target experiments at CERN due to the availability of a slow extraction from the SPS (multiple seconds long spill duration) at the long straight section 2 (LSS2) and of the existing infrastructure and beamlines. The former CERN Neutrinos to Gran Sasso (CNGS) experiment, however, was located deep underground downstream of LSS4 in TCC4 to be able to point the neutrino beam towards Gran Sasso, Italy. The ideal location in CERN’s North Area could be the ECN3 cavern due to its depth and radiation-protection limits. However, following the recent Research Board decision at CERN, the SHiP [36] experiment will occupy that cavern starting from LHC Run 4. The beamline cannot be placed inside the surface halls EHN1 and EHN2 in CERN’s North Area due to radiation constraints. The expected spill intensity exceeds and the secondary beam is of the order of , while the radiation protection (RP) constraints for EHN1 and EHN2 (supervised areas) are two to four orders of magnitude lower than these values even in a future EHN2 scenario with additional shielding that is currently proposed in the context of the AMBER experiment [37]. As a result, implementing such a beamline in EHN1 or EHN2 would be highly challenging without extensive shielding modifications, which would also affect the operation of other North Area beamlines. Additionally, the beamline’s bending angle makes it difficult to implement in experimental halls like EHN1 and EHN2, as these halls are predominantly long rather than wide. A placement below EHN1 is also excluded due to the technical galleries below the building and the challenges of excavating underneath an existing building in the vicinity of the Lion river. As a result, the existing ProtoDUNEs at the North Area’s Neutrino Platform (see Fig. 5) cannot be used in their current location for the SBN@CERN beamline, as the neutrino detector must be placed no more than from the end of the decay tunnel.111A beamline bypassing EHN1 to approach the ProtoDUNEs has been excluded due to the vicinity of the fence and the SPS beam rigidity that results in a large bending radius. Hence, building extensions and an experimental hall for the neutrino detectors would be necessary in the North Area.
@@ 2.3.1 Implementation at ECN4
The implementation of the SBN@CERN beamline has similar requirements compared to BDF/SHiP; however, a crucial difference is the required PoT (overall and annually) as well as the spill intensity. The envisaged spill intensity of is a factor of four smaller than the one envisaged by SHiP, which is . In addition, the SBN beamline would consume up to a maximum of 2 –, a value approximately ten times smaller than the BDF/SHiP requirement. This lower intensity allows for a reduced amount of shielding, cooling, and dilution measures that are required by BDF/SHiP, and offers greater flexibility in the location of the experimental hall. Simultaneously, key components of the initial ECN4 for the BDF/SHiP proposal are nearly identical in the case of SBN@CERN:
@@
Transfer tunnel: The SBN beamline also requires an extraction tunnel downstream of a junction cavern, which must also be constructed. The tunnel is comparable in length to that of the ECN4 proposal for BDF/SHiP; however, the entire beamline and decay tunnel would be placed fully underground once sufficient separation from the junction cavern is reached. Since the beamline only requires a shielding that is comparable to K12/NA62 due to a similar target and PoT requirements, the beamline should have sufficiently small transverse dimensions to achieve a manageable excavation process.
Target complex: Unlike BDF/SHiP, SBN does not require a complex target system due to its significantly lower intensity and different operational needs. The proposed graphite target is expected to have a simple design similar to that of CNGS and may not require replacement, as CNGS operated with a single graphite target throughout its lifetime. A beam dump for the primary protons that do not interact with the target is required, nevertheless. Here, a device like the XTAX currently used in the K12 beamline for the NA62 experiment is proposed, i.e., a long element with different layers of Al, Fe, and Cu.
@@ 2.3.2 Implementation at SPS LSS6
Apart from the North Area Extraction in LSS2, the two other long straight sections LSS4 and LSS6 (see Fig. 5) also feature extraction septa followed by transfer tunnels that serve the LHC Beam 2 and TCC4 (AWAKE) in the case of LSS4 and LHC Beam 1 and TCC6/TNC (HiRadMat) in the case of LSS6. These locations are deep underground and can be considered to house the proposed SBN@CERN facility. The main disadvantage compared to a location close to LSS2, however, is the missing ZS septa that are necessary for a slow extracted beam. Given the presence of AWAKE downstream in TCC4 and the depth of TCC4, placing the beamline downstream of LSS6 is considered more advantageous.
@@
If the non-local solution is pursued, some modifications of the SPS magnetic lattice have to be performed. A key aspect of this setup still has to be studied thoroughly – that is, the simultaneous extraction to LSS2 (North Area) and LSS6 (location of the SBN beamline). If accomplished, the simultaneous extraction would allow the beamline to be operated without competing for protons with other projects. Significant beam simulation and optics optimization studies have to be performed to ensure correct tune and resonance-driving term matching in the SPS before such an operation mode can be deemed feasible.
A potential placement of the SBN beamline in the TNC or TT61 tunnel has the advantage of the primary production target being deep underground. The TNC tunnel with a width of is slightly wider than the TT61 tunnel ( wide). The former TCC6 cavern (see Fig. 8; cavern upstream of TNC and TT61) housed the T1 target for the former West Area, receiving large numbers of protons on a yearly basis (more than ). HiRadMat also features a beam dump in the TNC tunnel; however, here it is worth noticing that HiRadMat can only receive due to RP constraints.
@@ 2.4 Proton sharing at the SPS
In the assumed scenario, the beamline is considered to be operated within a shared cycle with the other North Area experiments (SFTPRO cycle). This assumption is independent of the beamline placement, and a placement in TT61/TNC is also possible. A neutrino detector with a liquid argon mass is assumed to be located at a distance of from the end of the decay tunnel, requiring PoT to collect up to () charged-current (CC) events. The assumed SPS Flattop duration is s, delivering PoT per spill to the SBN beamline in a shared cycle. This corresponds to of the TCC2 intensity per spill and, ultimately, per year. ECN3 (BDF/SHiP) receives the maximum requested PoT/year. The SBN runtime for CC events in the 500 ton detector is then between and with a yearly PoT consumption of to . After five years, the statistical uncertainty on the inclusive cross-section would be better than in all cases.
In case the SPS Flattop duration would remain at the current value of (see [38]), the SBN beamline would still take roughly of the TCC2 intensity () and would reach the physics goal of within and . A statistical uncertainty of 1% on the inclusive cross-section is therefore likely to be achievable within five years. It is important to mention the possibility of an enhanced spill number for TCC2 in case the SHiP spill intensity is increased to . This would increase the number of spills for the North Area (SHiP excluded) by roughly and would shorten the experiment’s required runtime.
A second slow-extraction setup for extraction at either LSS4 or LSS6 of the SPS could provide the necessary flexibility to achieve simultaneous extractions to the SBN as well as the North Area (SFTPRO and SHiP cycles). This possibility is described in more detail in the corresponding PBC note, which analyzes a potential second slow-extraction setup in the SPS [35]. It would thus be possible to extract a small fraction of the beam during a SHiP cycle (at the percent level) in a simultaneous slow extraction. The cycle intensity can then be increased by that amount to still provide the target intensity to the SHiP experiment. This would allow for the reduction of the PoTs taken from the SFTPRO cycle. Such a more complicated sharing scenario will be studied further in the future. It is important to note that the beamline may never be operated in a dedicated cycle in either scenario. This is necessary to ensure coexistence with the other experiments while minimizing competition for PoTs.
@@ 3.2 Instrumented hadron dump
The low-energy component of the flux is produced by decays. This flux cannot be monitored through the instrumented tunnel because the muons produced in pairs with the neutrinos in pion decays are almost collinear with the axis of the tunnel, and thus outside the geometric acceptance of the calorimeter. A constraint of this flux component can only be achieved by installing detectors downstream of the hadron dump, allowing the measurement of forward muons exiting the tunnel. This places very stringent requirements on the detector technology and the optimisation of the detector layout. Nevertheless, the narrow-band beam and the slow proton extraction time of the SBN@CERN facility are useful for identifying these muons and achieving precise control of the neutrino flux.
@@ 4 Neutrino detectors
SBN@CERN aims to be the reference facility for neutrino cross-section measurements in the DUNE and Hyper-Kamiokande era. To maximize its potential, the facility requires detectors that can fully leverage the unprecedented precision of the beam while maintaining an optimal target for next-generation long-baseline cross-section studies. The two primary detectors will therefore utilize argon and water targets, eliminating systematic uncertainties associated with extrapolating nuclear effects across different atomic numbers (). The detailed detector choice will also depend on the site implementation of SBN@CERN, but the total (fiducial) detector mass is requested to be in the 500-1000 ton range, given the intensity and integrated protons of SBN@CERN.
@@ 4.1 Liquid-argon detector
The leading candidates for the argon-based detectors [55] at SBN@CERN are the ProtoDUNEs [56] and their future upgrades. During LHC Run 4, the existing ProtoDUNEs will undergo further upgrades to support the validation of the DUNE Phase II detectors [57]. Compared to the current ProtoDUNEs, all proposed technologies share a common set of characteristics. Most importantly, they preserve the imaging and particle reconstruction capabilities of DUNE at the GeV scale, ensuring that the primary objective of Phase II—precision observation of neutrino oscillations with the LBNF beam—remains fully achievable. A key advancement in the Phase-II detectors is the enhanced performance of the photon detection system, an upgrade also planned for DUNE’s third module, which is expected to be based on APEX or POWER technologies [57, 58]. This new photon detection system will play a crucial role in determining DUNE’s energy resolution at the few-MeV scale, significantly extending the experiment’s physics reach in solar and supernova neutrino studies while also providing additional sensitivity to BSM searches. The fourth module, often referred to as the “Module of Opportunity,” is still in the R&D phase. In addition to improved photon detection, it is expected to incorporate a pixelated charge readout, which would enhance spatial resolution, lower the energy threshold at the MeV scale, improve trigger efficiency by voxelizing the fiducial volume, and strengthen the identification of low-energy photons while improving radiogenic background rejection.
None of these improvements are mandatory to achieve the physics goals of SBN@CERN, as the DUNE Phase I detectors—and consequently the ProtoDUNEs—were optimized for optimal performance in the 0.5–5 GeV range. However, the capability to tag neutrinos beyond monitoring requires a time resolution better than in the neutrino detectors. Since the time resolution for GeV-scale neutrino events critically depends on the number of detected photoelectrons, enhancing the photon detection system (PDS) becomes essential to enable neutrino tagging in liquid argon.
As a result, the ideal detector for SBN@CERN would be one of the current ProtoDUNEs equipped with an upgraded timing system derived from the Phase II upgrades. Additionally, the implementation of a dedicated cosmic ray veto system to enhance cosmic background suppression is particularly valuable, given that SBN@CERN operates with a slow proton extraction. The design of this veto system, along with the potential installation of a muon catcher for full event containment, is currently under evaluation.

### v2

@@ 7 Neutrino cross-section measurements
- 7.8 Summary of the nuSCOPE physics potential
@@ 1 A high precision neutrino beam at CERN
In 2023, the convergence of the ENUBET and NuTag approaches revealed the potential to achieve neutrino-nucleus cross-section measurements with a precision comparable to electron-nucleus cross-sections. Recognizing this opportunity, the two collaborations joined forces to explore the feasibility of implementing such a facility at CERN [19]. The beamline design was optimized within the framework of Physics Beyond Colliders (PBC) at CERN, culminating in the design presented in this document. This document outlines the results of these efforts and provides the foundation for the proposal of a new high-precision, short-baseline neutrino beam at CERN, tentatively named SBN@CERN. In June 2025, an international collaboration was established, and the facility was officially named nuSCOPE – neutrino SPS COmplex for Precision Experiments. This facility aims to deliver neutrino cross-section measurements at the percent level in the energy range relevant for DUNE and Hyper-Kamiokande. While alternative implementations of this innovative technology outside CERN are under consideration, CERN’s unparalleled infrastructure, proton driver, and existing neutrino detectors position CERN as the most cost-effective venue for realizing this project in compliance with the CERN fixed-target program for the coming decade.
The nuSCOPE beamline is detailed in Sec. 2, with a focus on its implementation at CERN, utilizing the SPS as the proton driver. Section 3 outlines the beamline instrumentation, which monitors and tracks mesons and charged leptons, while Sec. 4 discusses the neutrino detectors that leverage the advancements of the CERN Neutrino Platform. Section 5 summarizes the technique that enables a sub-percent precision measurement of the neutrino flux using lepton monitoring and tracking. Section 6 presents the two methods by which nuSCOPE can measure the neutrino energy without relying on neutrino interactions: the Narrow Band Off-Axis technique developed by ENUBET and the SBND Collaboration at Fermilab [20, 21], which achieves a precision of approximately %, and the tagging method, which can reach the ultimate precision of 1%. The expected performance of the facility for a range of reference cross-section measurements is detailed in Sec. 7.
@@ 2 Beamline design
The nuSCOPE beamline aims at producing kaons and pions with a central momentum range of . The beamline, shown in Fig. 1, was optimized to produce (ideally) a monochromatic meson beam that remains parallel throughout a 40 m-long decay tunnel. The neutrino detectors that are under consideration are discussed in Sec. 4. In the following, it is assumed to be a liquid-argon detector located downstream of the decay tunnel. It has a transverse size of m2 and a longitudinal depth of roughly (see Sec. 4.3). The meson production is achieved via a thin graphite target that is impinged by the proton beam of CERN’s SPS.
@@
At CERN, the beamline can only be effectively driven by the SPS accelerator with a proton-beam energy in the range up to . The preliminary beamline specifications are listed in Tab. 1. At PS energies, the meson yield is highly suppressed making this option not viable. If implemented at CERN, the nuSCOPE beamline will be confronted with the issue of having a limited number of protons-on-target (PoT) available since other experiments are also dependent on a high annual proton flux from the SPS. For this reason, the beamline design has been optimized to minimize the required number of PoTs.
@@ 2.1 Beamline optimization
| Particle yield | ENUBET design | Optimized nuSCOPE design |
@@ 2.2 Radiation-protection considerations
nuSCOPE is a proposed beamline featuring novel instrumentation for flux monitoring and tagging. However, radiation effects must be carefully considered, as high radiation doses may impact beamline components, silicon taggers, and beamline instrumentation. This challenge has already been addressed by the ENUBET R&D and is further mitigated by the reduction of PoT achieved through the optimization described in Sec. 2.1. Concerning the ambient dose equivalent (H*(10)) studies are currently ongoing.
@@
Although radiation damage is not critical for the nuSCOPE components up to PoT, shielding for radiation protection within the CERN complex remains an important task. The required shielding depends on the final location of nuSCOPE and is currently under evaluation. Location options are discussed in the next section.
@@ 2.3 Possible implementation at CERN
The feasibility study for a potential placement of the nuSCOPE beamline on the CERN campus is currently ongoing and is not final [35]; however, two approaches have been identified as promising and could lead to the successful implementation of the beamline.
CERN’s North Area (see Fig. 5) is usually the prime location for fixed-target experiments at CERN due to the availability of a slow extraction from the SPS (multiple seconds long spill duration) at the long straight section 2 (LSS2) and of the existing infrastructure and beamlines. The former CERN Neutrinos to Gran Sasso (CNGS) experiment, however, was located deep underground downstream of LSS4 in TCC4 to be able to point the neutrino beam towards Gran Sasso, Italy. The ideal location in CERN’s North Area could be the ECN3 cavern due to its depth and radiation-protection limits. However, following the recent Research Board decision at CERN, the SHiP [36] experiment will occupy that cavern starting from LHC Run 4. The beamline cannot be placed inside the surface halls EHN1 and EHN2 in CERN’s North Area due to radiation constraints. The expected spill intensity exceeds and the secondary beam is of the order of , while the radiation protection (RP) constraints for EHN1 and EHN2 (supervised areas) are two to four orders of magnitude lower than these values even in a future EHN2 scenario with additional shielding that is currently proposed in the context of the AMBER experiment [37]. As a result, implementing such a beamline in EHN1 or EHN2 would be highly challenging without extensive shielding modifications, which would also affect the operation of other North Area beamlines. Additionally, the beamline’s bending angle makes it difficult to implement in experimental halls like EHN1 and EHN2, as these halls are predominantly long rather than wide. A placement below EHN1 is also excluded due to the technical galleries below the building and the challenges of excavating underneath an existing building in the vicinity of the Lion river. As a result, the existing ProtoDUNEs at the North Area’s Neutrino Platform (see Fig. 5) cannot be used in their current location for the nuSCOPE beamline, as the neutrino detector must be placed no more than from the end of the decay tunnel.111A beamline bypassing EHN1 to approach the ProtoDUNEs has been excluded due to the vicinity of the fence and the SPS beam rigidity that results in a large bending radius. Hence, building extensions and an experimental hall for the neutrino detectors would be necessary in the North Area.
@@ 2.3.1 Implementation at ECN4
The implementation of the nuSCOPE beamline has similar requirements compared to BDF/SHiP; however, a crucial difference is the required PoT (overall and annually) as well as the spill intensity. The envisaged spill intensity of is a factor of four smaller than the one envisaged by SHiP, which is . In addition, the nuSCOPE beamline would consume up to a maximum of 2 –, a value approximately ten times smaller than the BDF/SHiP requirement. This lower intensity allows for a reduced amount of shielding, cooling, and dilution measures that are required by BDF/SHiP, and offers greater flexibility in the location of the experimental hall. Simultaneously, key components of the initial ECN4 for the BDF/SHiP proposal are nearly identical in the case of nuSCOPE:
@@
Transfer tunnel: The nuSCOPE beamline also requires an extraction tunnel downstream of a junction cavern, which must also be constructed. The tunnel is comparable in length to that of the ECN4 proposal for BDF/SHiP; however, the entire beamline and decay tunnel would be placed fully underground once sufficient separation from the junction cavern is reached. Since the beamline only requires a shielding that is comparable to K12/NA62 due to a similar target and PoT requirements, the beamline should have sufficiently small transverse dimensions to achieve a manageable excavation process.
Target complex: Unlike BDF/SHiP, nuSCOPE does not require a complex target system due to its significantly lower intensity and different operational needs. The proposed graphite target is expected to have a simple design similar to that of CNGS and may not require replacement, as CNGS operated with a single graphite target throughout its lifetime. A beam dump for the primary protons that do not interact with the target is required, nevertheless. Here, a device like the XTAX currently used in the K12 beamline for the NA62 experiment is proposed, i.e., a long element with different layers of Al, Fe, and Cu.
@@ 2.3.2 Implementation at SPS LSS6
Apart from the North Area Extraction in LSS2, the two other long straight sections LSS4 and LSS6 (see Fig. 5) also feature extraction septa followed by transfer tunnels that serve the LHC Beam 2 and TCC4 (AWAKE) in the case of LSS4 and LHC Beam 1 and TCC6/TNC (HiRadMat) in the case of LSS6. These locations are deep underground and can be considered to house the proposed nuSCOPE facility. The main disadvantage compared to a location close to LSS2, however, is the missing ZS septa that are necessary for a slow extracted beam. Given the presence of AWAKE downstream in TCC4 and the depth of TCC4, placing the beamline downstream of LSS6 is considered more advantageous.
@@
If the non-local solution is pursued, some modifications of the SPS magnetic lattice have to be performed. A key aspect of this setup still has to be studied thoroughly – that is, the simultaneous extraction to LSS2 (North Area) and LSS6 (location of the nuSCOPE beamline). If accomplished, the simultaneous extraction would allow the beamline to be operated without competing for protons with other projects. Significant beam simulation and optics optimization studies have to be performed to ensure correct tune and resonance-driving term matching in the SPS before such an operation mode can be deemed feasible.
A potential placement of the nuSCOPE beamline in the TNC or TT61 tunnel has the advantage of the primary production target being deep underground. The TNC tunnel with a width of is slightly wider than the TT61 tunnel ( wide). The former TCC6 cavern (see Fig. 8; cavern upstream of TNC and TT61) housed the T1 target for the former West Area, receiving large numbers of protons on a yearly basis (more than ). HiRadMat also features a beam dump in the TNC tunnel; however, here it is worth noticing that HiRadMat can only receive due to RP constraints.
@@ 2.4 Proton sharing at the SPS
In the assumed scenario, the beamline is considered to be operated within a shared cycle with the other North Area experiments (SFTPRO cycle). This assumption is independent of the beamline placement, and a placement in TT61/TNC is also possible. A neutrino detector with a liquid argon mass is assumed to be located at a distance of from the end of the decay tunnel, requiring PoT to collect up to () charged-current (CC) events. The assumed SPS Flattop duration is s, delivering PoT per spill to the nuSCOPE beamline in a shared cycle. This corresponds to of the TCC2 intensity per spill and, ultimately, per year. ECN3 (BDF/SHiP) receives the maximum requested PoT/year. The nuSCOPE runtime for CC events in the 500 ton detector is then between and with a yearly PoT consumption of to . After five years, the statistical uncertainty on the inclusive cross-section would be better than in all cases.
In case the SPS Flattop duration would remain at the current value of (see [38]), the nuSCOPE beamline would still take roughly of the TCC2 intensity () and would reach the physics goal of within and . A statistical uncertainty of 1% on the inclusive cross-section is therefore likely to be achievable within five years. It is important to mention the possibility of an enhanced spill number for TCC2 in case the SHiP spill intensity is increased to . This would increase the number of spills for the North Area (SHiP excluded) by roughly and would shorten the experiment’s required runtime.
A second slow-extraction setup for extraction at either LSS4 or LSS6 of the SPS could provide the necessary flexibility to achieve simultaneous extractions to nuSCOPE as well as the North Area (SFTPRO and SHiP cycles). This possibility is described in more detail in the corresponding PBC note, which analyzes a potential second slow-extraction setup in the SPS [35]. It would thus be possible to extract a small fraction of the beam during a SHiP cycle (at the percent level) in a simultaneous slow extraction. The cycle intensity can then be increased by that amount to still provide the target intensity to the SHiP experiment. This would allow for the reduction of the PoTs taken from the SFTPRO cycle. Such a more complicated sharing scenario will be studied further in the future. It is important to note that the beamline may never be operated in a dedicated cycle in either scenario. This is necessary to ensure coexistence with the other experiments while minimizing competition for PoTs.
@@ 3.2 Instrumented hadron dump
The low-energy component of the flux is produced by decays. This flux cannot be monitored through the instrumented tunnel because the muons produced in pairs with the neutrinos in pion decays are almost collinear with the axis of the tunnel, and thus outside the geometric acceptance of the calorimeter. A constraint of this flux component can only be achieved by installing detectors downstream of the hadron dump, allowing the measurement of forward muons exiting the tunnel. This places very stringent requirements on the detector technology and the optimisation of the detector layout. Nevertheless, the narrow-band beam and the slow proton extraction time of the nuSCOPE facility are useful for identifying these muons and achieving precise control of the neutrino flux.
@@ 4 Neutrino detectors
nuSCOPE aims to be the reference facility for neutrino cross-section measurements in the DUNE and Hyper-Kamiokande era. To maximize its potential, the facility requires detectors that can fully leverage the unprecedented precision of the beam while maintaining an optimal target for next-generation long-baseline cross-section studies. The two primary detectors will therefore utilize argon and water targets, eliminating systematic uncertainties associated with extrapolating nuclear effects across different atomic numbers (). The detailed detector choice will also depend on the site implementation of nuSCOPE, but the total (fiducial) detector mass is requested to be in the 500-1000 ton range, given the intensity and integrated protons of nuSCOPE.
@@ 4.1 Liquid-argon detector
The leading candidates for the argon-based detectors [55] at nuSCOPE are the ProtoDUNEs [56] and their future upgrades. During LHC Run 4, the existing ProtoDUNEs will undergo further upgrades to support the validation of the DUNE Phase II detectors [57]. Compared to the current ProtoDUNEs, all proposed technologies share a common set of characteristics. Most importantly, they preserve the imaging and particle reconstruction capabilities of DUNE at the GeV scale, ensuring that the primary objective of Phase II—precision observation of neutrino oscillations with the LBNF beam—remains fully achievable. A key advancement in the Phase-II detectors is the enhanced performance of the photon detection system, an upgrade also planned for DUNE’s third module, which is expected to be based on APEX or POWER technologies [57, 58]. This new photon detection system will play a crucial role in determining DUNE’s energy resolution at the few-MeV scale, significantly extending the experiment’s physics reach in solar and supernova neutrino studies while also providing additional sensitivity to BSM searches. The fourth module, often referred to as the “Module of Opportunity,” is still in the R&D phase. In addition to improved photon detection, it is expected to incorporate a pixelated charge readout, which would enhance spatial resolution, lower the energy threshold at the MeV scale, improve trigger efficiency by voxelizing the fiducial volume, and strengthen the identification of low-energy photons while improving radiogenic background rejection.
None of these improvements are mandatory to achieve the physics goals of nuSCOPE, as the DUNE Phase I detectors—and consequently the ProtoDUNEs—were optimized for optimal performance in the 0.5–5 GeV range. However, the capability to tag neutrinos beyond monitoring requires a time resolution better than in the neutrino detectors. Since the time resolution for GeV-scale neutrino events critically depends on the number of detected photoelectrons, enhancing the photon detection system (PDS) becomes essential to enable neutrino tagging in liquid argon.
As a result, the ideal detector for nuSCOPE would be one of the current ProtoDUNEs equipped with an upgraded timing system derived from the Phase II upgrades. Additionally, the implementation of a dedicated cosmic ray veto system to enhance cosmic background suppression is particularly valuable, given that nuSCOPE operates with a slow proton extraction. The design of this veto system, along with the potential installation of a muon catcher for full event containment, is currently under evaluation.